How many grams of h20 will be produced by 6.30 moles of o2


1
Expert's answer
2021-03-03T02:01:27-0500

1 mole of O2 will produce 2 moles of H2O.

So, 6.30 moles of O2 will produce 2 × 6.30 = 12.60 moles of H2O.

Molar mass of H2O = 18 g/mol.

Therefore, the mass of H2O is produced = 12.60 × 18 = 226.80 grams.